### v1

@@ Abstract
Artificial Night-Time Light (NTL) remote sensing is a vital proxy for quantifying the intensity and spatial distribution of human activities. Although the NPP-VIIRS sensor provides high-quality NTL observations, its temporal coverage, which begins in 2012, restricts long-term time-series studies that extend to earlier periods. Despite the progress in extending VIIRS-like NTL time-series, current methods still suffer from two significant shortcomings: the underestimation of light intensity and the structural omission. To overcome these limitations, we propose a novel reconstruction framework consisting of a two-stage process: construction and refinement. The construction stage features a Hierarchical Fusion Decoder (HFD) designed to enhance the fidelity of the initial reconstruction. The refinement stage employs a Dual Feature Refiner (DFR), which leverages high-resolution impervious surface masks to guide and enhance fine-grained structural details. Based on this framework, we developed the Extended VIIRS-like Artificial Nighttime Light (EVAL) product for China, extending the standard data record backwards by 26 years to begin in 1986. Quantitative evaluation shows that EVAL significantly outperforms existing state-of-the-art products, boosting the $\text{R}^{2}$ from 0.68 to 0.80 while lowering the RMSE from 1.27 to 0.99. Furthermore, EVAL exhibits excellent temporal consistency and maintains a high correlation with socioeconomic parameters, confirming its reliability for long-term analysis. The resulting EVAL dataset provides a valuable new resource for the research community and is publicly available at https://doi.org/10.11888/HumanNat.tpdc.302930.
keywords:
[…]
1 Introduction
@@
Night-Time Light (NTL) data, as a unique means of Earth observation, precisely captures the radiant characteristics of light sources originating from nocturnal human activities, providing a critical dimension of information unattainable through conventional daytime remote sensingElvidge et al. (1997); Zheng et al. (2023); Levin et al. (2020). Unlike daytime remote sensing, which primarily relies on reflected solar radiation signals, NTL data directly records the intensity and distribution of light sources from human societal activities, establishing a direct link between human presence and environmental impact. With the accelerating pace of global urbanization and the expansion of human activities, NTL data have demonstrated significant scientific value in diverse fields such as land use change monitoringChen et al. (2022a); Ma et al. (2023); Zhang et al. (2023b); Zhang and Seto (2011), socioeconomic development assessmentBennett and Smith (2017); Shi et al. (2016); Sutton (1997); Ma et al. (2014), urbanization process analysisElvidge et al. (1997); Chen et al. (2022b); Shi et al. (2014); Henderson et al. (2003), carbon emission estimationJung et al. (2022); Fang et al. (2022); Xu et al. (2023), and ecological environment change monitoringMa et al. (2023); Zhang et al. (2023a); Gaston et al. (2013). This facilitates the exploration of complex interaction mechanisms between human activities and global change, providing a scientific basis for sustainable development goals.
@@
Currently, NTL remote sensing data are primarily sourced from two major satellite systems operated by the National Oceanic and Atmospheric Administration (NOAA): the Operational Linescan System (OLS) on the Defense Meteorological Satellite Program (DMSP), and the Visible Infrared Imaging Radiometer Suite (VIIRS) aboard the Suomi National Polar-orbiting Partnership (NPP). Specifically, DMSP-OLS NTL data cover a continuous observational record from 1992 to 2013, offering a relatively complete history of nighttime light variationsWu et al. (2021). In contrast, NPP-VIIRS NTL data, operational since 2012, boast superior spatial resolution and radiometric sensitivityShi et al. (2014). However, these two systems employ different measurement principles, radiometric calibration methods, and data quantization units, resulting in data formats and numerical ranges that lack direct comparability. This heterogeneity significantly limits their applicability in long-term time-series analyses. Without specialized cross-platform conversion, integrated use of these datasets is infeasible, thereby restricting the scope and depth of long-term studies based on NTL dataLevin et al. (2020).
@@
In response to these challenges, scholars in recent years have dedicated efforts to developing various data conversion and calibration methods to construct unified, long-term NTL datasets. Mainstream integration methods often use DMSP-OLS as the baselineWu et al. (2021); Zhang and Seto (2011); however, such approaches inherit the inherent deficiencies of the original DMSP-OLS dataLevin et al. (2020); Cao et al. (2019), such as low spatial resolution and over-saturation of pixels, which severely compromises the accuracy and reliability of the resulting data. In contrast, NPP-VIIRS data offer significant technical advantages in spatial resolution and radiometric precision, making them more suitable as a benchmark for long-term NTL data integration. However, research focused on constructing long-term NPP-VIIRS-like NTL data remains relatively scarce. In this frontier, Chen et al.Chen et al. (2020) proposed a cross-sensor calibration method, introducing the Enhanced Vegetation Index (EVI) as an auxiliary variable to augment the original DMSP NTL data into an Enhanced Artificial Nighttime Light Index (EANTLI), partially overcoming the issue of information loss in DMSP data. Subsequently, their research combined an autoencoder model to convert DMSP-OLS NTL data into an NPP-VIIRS-like format, generating a global NPP-VIIRS-like NTL dataset for the period of 2000 to 2018 named LongNTL. Building upon this, Chen et al.Chen et al. (2024) further refined the conversion method by integrating original DMSP-OLS NTL data with annual Normalized Difference Vegetation Index (NDVI) data. They introduced a U-Net deep learning model as the core architecture for the conversion, which preserves both global semantic information and local detail features, ultimately producing a global Simulated VIIRS Nightime Light Dataset (SVNL) covering the years 1992 to 2023.
@@
Although previous studies have established initial cross-sensor mappings from DMSP-OLS to NPP-VIIRS NTL data, enabling long-term integration, several technical limitations remain such as underestimation of light intensity and structural omission. First, existing methods underperform in regions with high radiance values—such as central business districts, industrial complexes, and ports—due to the saturation inherent in DMSP-OLS data. In these areas, the loss of gradient information severely hampers the model’s ability to reconstruct light intensity variations, resulting in distorted urban representations and biased estimates of socioeconomic parameters. Second, current VIIRS-like datasets omit intra-urban structures and road networks, limiting their effectiveness for analyzing regional development and transportation infrastructure. This shortcoming arises primarily from the auxiliary features employed—namely, spectral indices like EVI and NDVI—that are optimized for vegetated surfaces. These indices exhibit low sensitivity in non-vegetated areas such as urban built-up zones, industrial regions, transportation corridors, and arid landscapesSmall (2001), making it difficult to distinguish between anthropogenic and natural barren surfaces. As a result, these models struggle to reconstruct fine-scale urban structures and road connectivity. Moreover, most prior approaches rely on generic AutoencoderNg et al. (2011) or U-NetRonneberger et al. (2015) architectures. This methods are not specifically optimized for the unique challenges of NTL reconstruction, leaving considerable room for methodological improvement.
@@
In summary, although existing approaches for generating NPP-VIIRS-like NTL data have achieved notable progress, they continue to suffer from two key deficiencies: intensity underestimation and structural omission. These shortcomings limit their suitability for long-term time-series analyses. To overcome these challenges, this study introduces a novel framework specifically designed to produce high-quality NPP-VIIRS-like NTL data. Central to this framework is a specialized fine-tuning module appended to the backbone network. This module utilizes high-resolution impervious surface area masks to guide fine-grained corrections, thereby enhancing the structural fidelity and surface texture representation of the output data. Leveraging this methodology, we constructed an annual NPP-VIIRS-like NTL dataset for the years starting from 1986, which we name Extended VIIRS-like Artificial Nighttime Light (EVAL).
The main contributions of this study are summarized as follows:
  • We propose a new framework for reconstructing NPP-VIIRS-like NTL data based on DMSP NTL and multi-resolution auxiliary surface features. This framework consists of two sequential stages: a construction stage that performs initial reconstruction, and a refinement stage that applies fine-grained adjustments. This two-stage design presents a novel direction for the reconstruction of high-quality NPP-VIIRS-like NTL data, while also providing the flexibility to incorporate future improvements based on new backbone models.
  • We design two key modules, named Hierarchical Fusion Decoder (HFD) and Dual Feature Refiner (DFR). The HFD improves reconstruction quality by enhancing sensitivity to high-frequency spatial features and fusing multi-scale information. The DFR module utilizes high-resolution impervious surface masks to refine structural information, thereby enhancing the continuity and realism of the reconstructed spatial patterns.
  • Based on the proposed framework, we have developed EVAL, a high-quality, 500m resolution, long-term time-series NPP-VIIRS-like NTL dataset for China covering the period since 1986. This dataset effectively addresses the problems of underestimation and omission found in existing data and provides continuous and consistent data support for related research endeavors.
2 Datasets
@@
Harmonized Global Nighttime Light DatasetLi et al. (2020) Served as DMSP-OLS NTL features for training and inference (1992-2013). https://doi.org/10.6084/m9.figshare.9828827.v2
PANDA-ChinaZhang et al. (2024) Served as DMSP-OLS NTL features for early-period time-serious (1986-1991). https://data.tpdc.ac.cn/en/data/e755f1ba-9cd1-4e43-98ca-cd081b5a0b3e/
Annual VNL V2Elvidge et al. (2021) Served as the ground truth NPP-VIIRS NTL data for model training and accuracy assessment. https://eogdata.mines.edu/nighttime_light/annual/v20/
@@
Global Artificial Impervious AreaGong et al. (2020) Served as the impervious surface area mask to guide model training and inference. https://data-starcloud.pcl.ac.cn/zh/resource/13
LongNTLChen et al. (2020) Used as a benchmark dataset for accuracy assessment. https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/YGIVCD
SVNLChen et al. (2024) Used as a benchmark dataset for accuracy assessment. https://doi.org/10.6084/m9.figshare.22262545.v8
@@
2.1 Nighttime Light Data
DMSP-OLS NTL:
The original DMSP-OLS NTL data suffer from temporal inconsistencies and fluctuations in data quality due to several issues, including the lack of on-board calibration, inter-sensor performance discrepancies, and sensor degradation over time. To overcome these limitations, this study utilizes the systematically, stepwise-calibrated global DMSP-OLS NTL dataset developed by Li et al.Li et al. (2020). This dataset exhibits enhanced temporal consistency and comparability. Consequently, it was selected as the baseline DMSP-like data for the period of 1992 to 2013, serving as the foundation for both the training and inference processes of our model.
To further extend the temporal scope of the research and cover periods where DMSP data is unavailable, this study incorporates the PANDA-China datasetZhang et al. (2024) as a supplement. In this dataset, we provide a high-precision and temporally consistent DMSP-like NTL data product spanning from 1984 to 2020. In this research, the PANDA-China dataset was employed for model inference during two specific intervals where the original DMSP data is absent: 1986-1991 and 2014-2020.
Considering the inherent dissimilarities between the PANDA-China and Li’s NTL datasets, we trained a simple U-Net modelRonneberger et al. (2015) for two separate years, 1992 and 2013. These models were respectively used to map the PANDA-China NTL data to the characteristics of Li’s NTL data for the 1986-1991 and 2014-2020 periods.
NPP-VIIRS NTL:
This study utilizes the annual mean composite NTL dataset from the NPP-VIIRS Day/Night BandElvidge et al. (2021) for the years 2012 to 2020, released by the Earth Observation Group at the Colorado School of Mines. This dataset is generated from monthly NPP-VIIRS data, during which interferences from ephemeral or anomalous light signals—such as those from fires, auroras, high-energy particle impacts, and background noise—are systematically removed in the synthesis process. To address the issue of missing data in parts of Northern China for several months in 2012, this study adheres to the outlier removal and annual mean calculation methodology proposed by Elvidge et al.Elvidge et al. (2021) to ensure consistency between the data for that year and subsequent years.
The research proceeds on the assumption that major urban centers exhibit the highest NTL intensity. A pixel is identified as an anomaly and adjusted if its NTL intensity surpasses the maximum observed in major cities. In the course of this study, four of the most developed representative cities in China were selected: Beijing, Shanghai, Hong Kong, and Taipei. The average of the maximum NTL intensity values from these four cities was established as a threshold. Any NPP-VIIRS NTL value exceeding this threshold was replaced with the average NTL intensity of its 24 neighboring pixels. According to the findings of Zhao et al.Zhao et al. (2019), the distribution of NPP-VIIRS NTL data after a logarithmic transformation more closely resembles that of DMSP-OLS NTL data. Therefore, during the model training and prediction phases, we used the natural log-transformed NPP-VIIRS NTL data as the target variable. Upon completion of the prediction, an exponential transformation was applied to map the generated results back to the original data space, thereby ensuring the output is consistent with the actual data.
2.2 Auxiliary Data
Landsat Surface Reflectance:
To acquire fine-scale land cover information, this study utilized the Tier 1 Surface Reflectance (SR) products from the Landsat series of satellites (TM, ETM+, and OLI)Arvidson et al. (2001); Wulder et al. (2019). The data were sourced from the United States Geological Survey (USGS) and included six spectral bandsLoveland and Irons (2016): Blue, Green, Red, Near-Infrared (NIR), Shortwave Infrared 1 (SWIR1), and Shortwave Infrared 2 (SWIR2). These SR products have undergone atmospheric correction using the LEDAPSMasek et al. (2006) and LaSRCVermote et al. (2016) algorithms employed by the USGS, which mitigates the impact of atmospheric effects on the surface reflectance signal. To further minimize the influence of clouds, cloud shadows, and snow/ice cover, this study computed the 10th percentile of the surface reflectance values for each spectral band from all available observations within a given year, which was then used as the representative annual surface reflectance. Furthermore, to ensure temporal consistency across data from different sensors, the SR data from the OLI sensor were normalized to the level of the TM and ETM+ sensors. This harmonization was based on the transformation coefficients proposed by Roy et al.Roy et al. (2016), thereby guaranteeing the comparability of the multi-temporal Landsat imagery across different years.
Artificial Impervious Surface:
To accurately capture the structure of the built environment within cities and the connectivity features between regions, this study incorporated the Global Artificial Impervious Areas (GAIA) datasetGong et al. (2020). In the GAIA dataset, we provide annual 30-meter spatial resolution masks of the Earth’s impervious surfaces covering the entire study period. This dataset is characterized by a global overall accuracy exceeding 90% and exhibits good time-series consistency.
3 Methodology
@@
3.1 Pipeline
The overall architecture of our proposed framework is illustrated in Figure. 1 and consists of a backbone network followed by an attached refiner module. The backbone model adopts a U-NetRonneberger et al. (2015) architecture, comprising an encoder, a decoder, and skip connections. It accepts DMSP NTL and surface reflectance data as input to reconstruct the log-transformed NPP-VIIRS NTL data.
Figure 1: The architecture of our proposed framework, detailing the overall pipeline and its key components. (a) The overall pipeline consists of a construction stage and a refinement stage. The construction stage uses a U-Net-based architecture to generate an initial prediction via our specialized decoder (HFD). The subsequent refinement stage then performs fine-grained adjustments on this prediction using the refiner module (DFR). (b) The detailed structure of the Hierarchical Fusion (HFD) decoder. It is composed of a Structure Residual Fusion (SRF) module, which intelligently incorporates fine-grained details from the encoder’s skip connection, and a Multi-scale Aggregator (MA), which captures and adaptively fuses contextual features from varying receptive fields. (c) The detailed structure of the Dual Feature Refiner (DFR) module. It utilizes residual blocks and a Cross-Resolution Local Attention (CLA) to fine-tune the reconstructed image, guided by high-resolution features.
Within the backbone a traditional visual encoder (e.g. ResNet-50He et al. (2016)), can be flexibly employed. It generates Multi-Scale feature maps through five successive downsampling stages. In the decoder section, we have designed five corresponding upsampling stages. Each stage takes as input the feature map from the preceding stage, the feature map from the corresponding encoder stage via skip connection. To address the specific task of nighttime light reconstruction, we have integrated a Structure Residual Fusion (SRF) module and a Mutiscale Aggregator (MA) within the decoder, aimed at enhancing the clarity of the reconstructed images.
Following the backbone network, we separately train the DFR. This module utilizes a Cross-Resolution Local Attention (CLA), leveraging high-resolution impervious surface area masks as guidance to optimize the expression of fine-grained structural features, which greatly improves the results for internal urban fabric and road networks.
@@ 3.1 Pipeline
3.2 Hierarchical Fusion Decoder
As illustrated in Figure 1 (b), the Hierarchical Fusion Decoder (HFD) takes as input the upsampled features from the preceding decoder layer and the skip-connection features from the corresponding encoder layer. The HFD is consists of SRF and MA.
Within the SRF module, we first process the skip-connection features through two parallel pathways. A 1×1 convolution extracts "content" information, while a 5×5 convolution extracts "structure" information. The main pathway combines the "content" information with the feature map upsampled by PixelShuffle via concatenation. Meanwhile, the "structure" information passes through a gating module to create an optimized structural residual. Finally, we add this residual to the main pathway. This design enhances the recovery of edges and textures while preserving high-level semantic information. While the MA is Inspired by DarkIRFeijoo et al. (2025). We design a parallel multi-branch structure following the initial fusion to capture rich contextual information. This structure consists of three parallel branches employing dilated convolutions with dilation rates of 1, 4, and 9. This configuration allows the model to obtain a wide receptive field, capturing features from local to near-global scales. To adaptively aggregate these Multi-Scale features, we introduce a lightweight channel attention mechanism that dynamically generates a distinct weight for each path. The final output of the module is a weighted sum of the features from each path, combined according to their learned weights.
3.3 Dual Feature Refiner
Following the backbone network, we introduce a cross-resolution refinement module, termed the Dual Feature Refiner, to perform localized, fine-grained corrections on the initial prediction. Panel (c) of Figure 1 illustrates the detailed structure of the DFR. The core idea of this module is to leverage high-resolution impervious surface area masks to guide the detailed fine-tuning of the low-resolution reconstructed image. Its final output is a residual correction which is then added to the preliminary low-resolution output to optimizes the final result while minimizing alterations to the initial prediction from the backbone network.
The Refiner employs a dual-branch, cross-resolution architecture. The low-resolution and high-resolution branches extract their respective deep features through a series of stacked residual blocks. Central to the fusion of information between the two branches is our proposed Cross-Resolution Local Attention (CLA). This mechanism discards computationally expensive global attention. Instead, it enables each pixel in the low-resolution feature map to attend to a corresponding 5x5 local neighborhood within the high-resolution feature map. Within this localized window, the module executes standard scaled dot-product attention. This allows the low-resolution features to adaptively extract the most relevant texture and edge details from the high-resolution reference information.
The low-resolution features, thus enhanced by the CLA, are subsequently fused with the downsampled high-resolution features. This fused representation is then passed through several additional residual blocks for final information integration before being used to predict the residual correction for the initial prediction.
3.4 Evaluation metrics

### v2

@@ Abstract
Artificial Night-Time Light (NTL) remote sensing is a vital proxy for quantifying the intensity and spatial distribution of human activities. Although the NPP-VIIRS sensor provides high-quality NTL observations, its temporal coverage, which begins in 2012, restricts long-term time-series studies that extend to earlier periods. Current extended VIIRS-like NTL data products suffer from two significant shortcomings: the underestimation of light intensity and the omission of structural details. To overcome these limitations, we present the Extended VIIRS-like Artificial Nighttime Light (EVAL) dataset, a new annual NTL dataset for China spanning from 1986 to 2024. This dataset was generated using a novel two-stage deep learning model designed to address the aforementioned shortcomings. The model first constructs an initial estimate and subsequently refines fine-grained structural details using high-resolution impervious surface data as guidance. Quantitative evaluations demonstrate that EVAL significantly outperforms state-of-the-art products, exhibiting superior temporal consistency and a stronger correlation with socioeconomic indicators.
Background & Summary
@@
Remote sensing of Night-Time Light (NTL) offers a unique modality for Earth observation by capturing the radiant characteristics of light sources originating from nocturnal human activities. This provides a critical dimension of information unattainable through conventional daytime remote sensing (elvidge1997mapping, ; zheng2023nighttime, ; levin2020remote, ). Unlike daytime remote sensing, which primarily relies on reflected solar radiation, NTL data directly record the intensity and distribution of artificial light, establishing a strong proxy for human presence and socioeconomic dynamics. With the accelerating pace of global urbanization and the expansion of human activities, NTL data have demonstrated significant scientific value in diverse fields such as land use change monitoring (chen2022global, ; ma2023global, ; zhang2011mapping, ), socioeconomic development assessment (bennett2017advances, ; sutton1997modeling, ; ma2014responses, ), urbanization process analysis (elvidge1997mapping, ; henderson2003validation, ; liu2024enhancing, ), carbon emission estimation (jung2022does, ; fang2022drives, ; xu2023spatio, ), and ecological environment change monitoring (ma2023global, ; zhang2023spatial, ; gaston2013ecological, ). Furthermore, NTL remote sensing plays a crucial role in advancing the United Nations Sustainable Development Goals (SDGs). It not only enables the direct assessment of progress towards SDG 7.1 (Affordable and Clean Energy) and SDG 11.1 (Sustainable Cities and Communities)(lv2025advancing, ; levin2025challenges, ), but also indirectly sheds light on other human activity-related goals, including SDG 1.1, SDG 1.2, SDG 8.1, and SDG 10.2(zhang2023spatial, ).
@@
Currently, NTL remote sensing data are primarily sourced from two major satellite systems: the Operational Linescan System (OLS) on the Defense Meteorological Satellite Program (DMSP), and the Visible Infrared Imaging Radiometer Suite (VIIRS) aboard the Suomi National Polar-orbiting Partnership (NPP). Specifically, DMSP-OLS NTL data cover a continuous observational record from 1992 to 2013, offering a relatively complete history of nighttime light variations(wu2021developing, ). In contrast, NPP-VIIRS NTL data, operational since 2012, boast superior spatial resolution and radiometric sensitivity(shi2014evaluation, ). However, these two systems employ different measurement principles, radiometric calibration methods, and data quantization units, resulting in data formats and numerical ranges that lack direct comparability. This heterogeneity significantly limits their applicability in long-term time-series analyses. Without specialized cross-platform conversion, integrated use of these datasets is infeasible, thereby restricting the scope and depth of long-term studies based on NTL data(levin2020remote, ).
@@
In response to these challenges, scholars in recent years have dedicated efforts to developing various data conversion and calibration methods to construct unified, long-term NTL datasets. Mainstream integration methods often use DMSP-OLS as the baseline. Zhang et al. (zhang2016robust, ) used ridge regression for calibration, while Li et al. (li2020harmonized, ) applied kernel density estimation to convert NPP-VIIRS data into a DMSP-OLS-like product. Similarly, Wu et al. (wu2021developing, ) adopted a pseudo-invariant pixel method. however, such approaches inherit the inherent deficiencies of the original DMSP-OLS data(levin2020remote, ), such as coarse spatial resolution and pixel over-saturation, which severely compromises the accuracy and reliability of the resulting data. Subsequent research targeted these specific issues: Cao et al. (cao2019simple, ) utilized a self-adjusting model to mitigate over-saturation, while Yu et al. (yu2025reconnecting, ) developed the 500-meter Tongji-NTL dataset through a super-resolution framework. Despite these improvements, a critical limitation persists: these datasets are not calibrated to the absolute radiometric units of NPP-VIIRS, which constrains their physical interpretation and quantitative use. This gap points toward a more robust paradigm: upgrading historical DMSP data to match the superior standard of NPP-VIIRS, which excels in both spatial detail and radiometric precision. In this frontier, Chen et al.(chen2020extended, ) proposed a cross-sensor calibration method, introducing the Enhanced Vegetation Index (EVI) as an auxiliary variable to augment the original DMSP NTL data into an Enhanced Artificial Nighttime Light Index (EANTLI), partially overcoming the issue of information loss in DMSP data. Subsequently, their research combined an autoencoder model to convert DMSP-OLS NTL data into an NPP-VIIRS-like format, generating a global NPP-VIIRS-like NTL dataset for the period of 2000 to 2018 named LongNTL. Building upon this, Chen et al.(chen2024global, ) further refined the conversion method by integrating original DMSP-OLS NTL data with annual Normalized Difference Vegetation Index (NDVI) data. They introduced a U-Net deep learning model as the core architecture for the conversion, which preserves both global semantic information and local detail features, ultimately producing a global Simulated VIIRS Nightime Light Dataset (SVNL) covering the years 1992 to 2023.
@@
Although previous studies have established initial cross-sensor mappings from DMSP-OLS to NPP-VIIRS NTL data, enabling long-term integration, several technical limitations remain such as underestimation of light intensity and structural omission. First, existing methods underperform in regions with high radiance values—such as central business districts, industrial complexes, and ports—due to the saturation inherent in DMSP-OLS data. In these areas, the loss of gradient information severely hampers the model’s ability to reconstruct light intensity variations, resulting in distorted urban representations and biased estimates of socioeconomic indicators. Second, current VIIRS-like datasets omit intra-urban structures and road networks, limiting their effectiveness for analyzing regional development and transportation infrastructure. This shortcoming arises primarily from the auxiliary features employed—namely, spectral indices like EVI and NDVI—that are optimized for vegetated surfaces. These indices exhibit low sensitivity in non-vegetated areas such as urban built-up zones, industrial regions, transportation corridors, and arid landscapes(small2001estimation, ), making it difficult to distinguish between anthropogenic and natural barren surfaces. As a result, these models struggle to reconstruct fine-scale urban structures and road connectivity. Moreover, most prior approaches rely on generic Autoencoder(ng2011sparse, ) or U-Net(ronneberger2015u, ). These methods primarily employ generic architectures that are not optimized for the unique challenges of NTL data, such as its extreme dynamic range and pixel saturation artifacts. Unlike standard daytime optical imagery, NTL data requires specialized handling to resolve these issues. Consequently, current models often struggle to accurately reconstruct high light intensities and fine spatial details.
@@
In summary, although existing approaches for generating NPP-VIIRS-like NTL data have achieved notable progress, they continue to suffer from two key deficiencies: intensity underestimation and structural omission. These shortcomings limit their suitability for long-term time-series analyses. To overcome these challenges, this study introduces a novel two-stage deep learning framework specifically designed to produce high-quality NPP-VIIRS-like NTL data. The framework first performs an initial reconstruction and subsequently refines fine-grained structural details by fusing multi-scale information and leveraging high-resolution impervious surface masks for guidance. Using this methodology, we constructed the Extended VIIRS-like Artificial Nighttime Light (EVAL) dataset. It provides a high-quality, annual, 500m resolution, NPP-VIIRS-like NTL time series for China, spanning the period from 1986 to 2024. This dataset effectively addresses the problems of underestimation and structural omission found in existing data and provides continuous and consistent data support for related research endeavors.
Methods
Data sources
@@ Methods
DMSP-OLS NTL
The original DMSP-OLS NTL data suffer from temporal inconsistencies and fluctuations in data quality due to several issues, including the lack of on-board calibration, inter-sensor performance discrepancies, and sensor degradation over time. To overcome these limitations, this study utilizes the systematically, stepwise-calibrated global DMSP-OLS NTL dataset developed by Li et al.(li2020harmoniz, ) (https://doi.org/10.6084/m9.figshare.9828827.v2). This dataset exhibits enhanced temporal consistency and comparability. Consequently, it was selected as the baseline DMSP-like data for the period of 1992 to 2013, serving as the foundation for both the training and inference processes of our model.
To further extend the temporal scope of the research and cover periods where DMSP data is unavailable, this study incorporates the PANDA-China dataset(zhang2024prolonged, ) as a supplement (https://data.tpdc.ac.cn/en/data/e755f1ba-9cd1-4e43-98ca-cd081b5a0b3e/). In this dataset, PANDA-China provide a high-precision and temporally consistent DMSP-like NTL data product spanning from 1984 to 2020. In this research, the PANDA-China dataset was employed for model inference during 1986-1991, the specific intervals where the original DMSP data is absent..
Considering the inherent dissimilarities between the PANDA-China and Li’s NTL datasets, we trained a simple U-Net model(ronneberger2015u, ) for the year 1992. This model was used to map the PANDA-China NTL data to the characteristics of Li’s NTL data for the 1986-1991period.
NPP-VIIRS NTL
This study utilizes the annual mean composite NTL dataset from the NPP-VIIRS Day/Night Band(elvidge2021annual, ) for the years 2012 to 2024 (https://eogdata.mines.edu/nighttime_light/annual/v20/), released by the Earth Observation Group at the Colorado School of Mines. This dataset is generated from monthly NPP-VIIRS data, during which interferences from ephemeral or anomalous light signals—such as those from fires, auroras, high-energy particle impacts, and background noise—are systematically removed in the synthesis process. To address the issue of missing data in parts of Northern China for several months in 2012, this study adheres to the outlier removal and annual mean calculation methodology proposed by Elvidge et al.(elvidge2021annual, ) to ensure consistency between the data for that year and subsequent years.
The research proceeds on the assumption that major urban centers exhibit the highest NTL intensity. A pixel is identified as an anomaly and adjusted if its NTL intensity surpasses the maximum observed in major cities. In the course of this study, four of the most developed representative cities in China were selected: Beijing, Shanghai, Hong Kong, and Taipei. The average of the maximum NTL intensity values from these four cities was established as a threshold. Any NPP-VIIRS NTL value exceeding this threshold was replaced with the average NTL intensity of its 24 neighboring pixels. According to the findings of Zhao et al.(zhao2019building, ), the distribution of NPP-VIIRS NTL data after a logarithmic transformation more closely resembles that of DMSP-OLS NTL data. Therefore, during the model training and prediction phases, we used the natural log-transformed NPP-VIIRS NTL data as the target variable. Upon completion of the prediction, an exponential transformation was applied to map the generated results back to the original data space, thereby ensuring the output is consistent with the actual data.
Landsat surface reflectance
To acquire fine-scale land cover information, this study utilized the Tier 1 Surface Reflectance (SR) products from the Landsat series of satellites (TM, ETM+, and OLI)(arvidson2001landsat, ; wulder2019current, ). The data were sourced from the United States Geological Survey (USGS) and included six spectral bands(loveland2016landsat, ): Blue, Green, Red, Near-Infrared (NIR), Shortwave Infrared 1 (SWIR1), and Shortwave Infrared 2 (SWIR2). These SR products have undergone atmospheric correction using the LEDAPS(masek2006landsat, ) and LaSRC(vermote2016preliminary, ) algorithms employed by the USGS, which mitigates the impact of atmospheric effects on the surface reflectance signal. To further minimize the influence of clouds, cloud shadows, and snow/ice cover, this study computed the 10th percentile of the surface reflectance values for each spectral band from all available observations within a given year, which was then used as the representative annual surface reflectance. Furthermore, to ensure temporal consistency across data from different sensors, the SR data from the OLI sensor were normalized to the level of the TM and ETM+ sensors. This harmonization was based on the transformation coefficients proposed by Roy et al.(roy2016characterization, ), thereby guaranteeing the comparability of the multi-temporal Landsat imagery across different years.
Artificial impervious surface
To accurately capture the structure of the built environment within cities and the connectivity features between regions, this study incorporated the Global Artificial Impervious Areas (GAIA) dataset(gong2020annual, ) (https://data-starcloud.pcl.ac.cn/zh/resource/13). GAIA provides annual 30-meter spatial resolution masks of the Earth’s impervious surfaces covering the entire study period. This dataset is characterized by a global overall accuracy exceeding 90% and exhibits good time-series consistency.
@@
Li’s Nighttime Light Dataset(li2020harmonized, ) Served as DMSP-OLS NTL features for training and inference (1992-2013). https://doi.org/10.6084/m9.figshare.9828827.v2
PANDA-China(zhang2021prolonged, ) Served as DMSP-OLS NTL features for early-period time-serious (1986-1991). https://doi.org/10.11888/Socioeco.tpdc.271202
Annual VNL V2(elvidge2021annual, ) Served as the ground truth NPP-VIIRS NTL data for model training and accuracy assessment. https://eogdata.mines.edu/nighttime_light/annual/v20/
@@
Global Artificial Impervious Area(gong2020annual, ) Served as the impervious surface area mask to guide model training and inference. https://data-starcloud.pcl.ac.cn/zh/resource/13
LongNTL(chen2020global, ) Used as a benchmark dataset for accuracy assessment. https://doi.org/10.7910/DVN/YGIVCD
SVNL(chen2023history, ) Used as a benchmark dataset for accuracy assessment. https://doi.org/10.6084/m9.figshare.22262545.v8
@@
Framework
@@
Pipeline
The overall architecture of our proposed framework is illustrated in Figure. 1 and consists of a backbone network followed by an attached refiner module. The backbone model adopts a U-Net(ronneberger2015u, ) architecture, comprising an encoder, a decoder, and skip connections. It accepts DMSP NTL and surface reflectance data as input to reconstruct the log-transformed NPP-VIIRS NTL data.
Figure 1: The architecture of our proposed framework, detailing the overall pipeline and its key components. (a) The overall pipeline consists of a construction stage and a refinement stage. The construction stage uses a U-Net-based architecture to generate an initial prediction via our specialized decoder (HFD). The subsequent refinement stage then performs fine-grained adjustments on this prediction using the refiner module (DFR). (b) The detailed structure of the HFD. It is composed of a Structure Residual Fusion (SRF) module, which intelligently incorporates fine-grained details from the encoder’s skip connection, and a Multi-scale Aggregator (MA), which captures and adaptively fuses contextual features from varying receptive fields. (c) The detailed structure of the DFR module. It utilizes residual blocks and a Cross-Resolution Local Attention to fine-tune the reconstructed image, guided by high-resolution features.
Within the backbone, a standard visual encoder produces multi-scale features through five downsampling stages, mirrored by five upsampling stages in the decoder that fuse encoder and decoder features via skip connections. To better address the challenges of nighttime light reconstruction, we introduce the HFD comprising a Structure Residual Fusion (SRF) module and a Multiscale Aggregator (MA). The SRF enhances structural detail and edge sharpness by refining skip features while suppressing mis-registered noise. The MA, in parallel, adaptively fuses multi-scale contextual cues to reduce over-spread artifacts. Together, they yield clearer and more consistent reconstructions.
Following the backbone network, we separately train the DFR. This module utilizes a Cross-Resolution Local Attention, leveraging high-resolution impervious surface area masks as guidance to optimize the expression of fine-grained structural features, which greatly improves the results for internal urban fabric and road networks.
@@
Hierarchical fusion decoder
As shown in Figure 1 (b), the HFD processes skip-connection features and upsampled features from the previous layer, and is composed of an SRF and an MA module. The SRF module processes skip-connections in parallel: a 1x1 convolution for ”content” and a 5x5 convolution for ”structure.” The ”content” is concatenated with the upsampled features, while the ”structure” information is added as a residual, enhancing edge and texture recovery. The MA module, inspired by DarkIR(feijoo2025darkir, ), then uses three parallel dilated convolution branches to capture multi-scale contextual information. A lightweight channel attention mechanism adaptively weights and sums the features from these branches.
Dual feature refiner
The DFR, illustrated in Figure 1 (c), is a cross-resolution module that performs localized, fine-grained corrections. It leverages high-resolution impervious surface area masks to guide the refinement of the initial NTL prediction by outputting a residual correction. The DFR employs a dual-branch architecture using residual blocks to extract features. Information is fused via a Cross-Resolution Local Attention mechanism, where each low-resolution pixel attends to a 5x5 local neighborhood in the high-resolution feature map using scaled dot-product attention. This allows for the adaptive extraction of relevant texture details. The enhanced low-resolution features are then fused with downsampled high-resolution features and passed through final residual blocks to predict the residual correction.
Evaluation metrics